OpenStudy (anonymous):

Zane bought a square sheet of paper of side 16 inches to decorate a pencil box he made in craft class. The pencil box has a height of 6 inches, a width of 3 inches, and a length of 5 inches. Approximately how many square inches of paper will be remaining if he pastes paper on to the bottom base and the four side surfaces of the pencil box without any overlap? Answer 130 in2 145 in2 160 in2 193 in2

OpenStudy (anonymous):

well since the SA of a rectangle prism is = 2lw + 2wh + 2lh, but since in this case it would only be one base it will look like this because it will take out on of the (lw). So its SA = lw + 2wh + 2lh SA = 5(3) + 2(3)(6) + 2(5)(6) SA = 15 + 36 + 60 SA = 111 Now the problem explains the square sheet so 16 * 16 would equal the SA of that paper. SA = 16 * 16 SA = 256 So we subtract the area of the pencil box she wants to cover from the papers area to find out the remaining paper. 256 - 111 = Remaining paper Remaining paper = 145 in^2

OpenStudy (anonymous):

thank-you!!